# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ 4.1.2 General AR/ARMA Setting
Table 1: DGPs (A) and (B). CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO. For each $n$ and $\phi$ setting the best CoEr and RMSFE are in bold.
DGP (A) (B) n 50 150 300 50 150 300 $\phi$ 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 SNR 1 CoEr LASSOy 1.00 0.99 0.98 0.98 1.00 1.00 0.99 0.99 1.00 1.00 1.00 0.99 0.99 0.86 0.54 0.56 0.99 0.88 0.66 0.69 0.99 0.90 0.73 0.75 GLS-LAS 0.94 0.75 0.76 0.81 0.93 0.76 0.78 0.72 0.94 0.79 0.83 0.75 0.96 0.81 0.64 0.72 0.96 0.82 0.82 0.85 0.97 0.84 0.87 0.88 ARDL-LAS 1.00 0.99 0.84 0.85 1.01 0.99 0.96 0.97 1.01 1.00 0.97 0.98 0.97 0.84 0.46 0.45 0.98 0.87 0.58 0.61 1.07 0.89 0.65 0.67 FaSel 3.92 2.75 1.48 1.39 2.02 1.37 1.00 0.89 1.38 1.10 0.94 0.86 1.04 0.94 1.03 1.02 0.98 0.91 1.05 1.06 3.93 2.30 1.05 1.06 ARMAr-LAS 0.96 0.73 0.45 0.44 0.96 0.75 0.54 0.47 0.96 0.77 0.56 0.48 0.98 0.81 0.41 0.39 0.98 0.82 0.52 0.53 0.99 0.83 0.58 0.58 RMSFE LASSOy 1.00 0.99 0.98 0.99 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.98 0.90 0.79 0.82 0.99 0.92 0.86 0.85 0.99 0.92 0.87 0.86 GLS-LAS 0.94 0.83 0.80 0.82 0.94 0.83 0.77 0.69 0.95 0.84 0.82 0.74 0.95 0.82 0.79 0.81 0.97 0.85 0.89 0.88 0.96 0.86 0.92 0.90 ARDL-LAS 1.01 1.01 1.00 1.01 1.01 1.01 1.00 0.99 1.01 1.01 0.99 1.01 0.99 0.89 0.72 0.73 0.99 0.90 0.82 0.79 1.02 0.91 0.83 0.80 FaSel 1.04 1.00 0.94 0.95 1.02 1.00 0.91 0.86 0.97 0.96 0.89 0.87 1.01 1.01 1.00 0.98 1.00 1.02 0.96 0.96 1.21 1.09 0.97 0.95 ARMAr-LAS 0.97 0.83 0.63 0.64 0.98 0.85 0.65 0.57 1.01 0.84 0.69 0.61 0.96 0.81 0.63 0.64 0.98 0.85 0.74 0.70 0.97 0.85 0.74 0.71 % TP LASSO 100.00 99.98 99.61 99.53 100.00 100.00 99.54 98.45 99.99 100.00 99.60 98.48 58.62 49.69 56.63 59.77 47.08 40.76 44.26 44.61 41.00 33.20 37.21 38.42 LASSOy 100.00 99.98 99.59 99.54 100.00 100.00 99.50 98.33 99.99 100.00 99.50 98.31 59.28 51.31 42.77 46.62 47.31 41.66 34.78 36.41 41.24 34.70 29.64 31.17 GLS-LAS 100.00 100.00 99.82 99.80 100.00 100.00 99.88 99.77 100.00 100.00 99.85 99.64 60.31 59.94 51.43 54.12 48.90 49.02 42.08 42.44 43.07 41.27 36.86 37.92 ARDL-LAS 100.00 99.98 99.50 99.41 100.00 100.00 99.50 98.24 99.99 100.00 99.44 98.15 57.89 52.08 48.15 51.23 46.51 42.38 37.70 38.95 41.61 35.25 30.99 33.08 FaSel 54.41 68.56 87.93 91.41 88.30 96.64 99.63 99.41 98.90 99.88 99.82 99.41 7.63 16.13 51.21 53.88 7.11 14.45 40.60 42.59 62.08 44.41 34.88 38.26 ARMAr-LAS 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.99 100.00 100.00 100.00 61.82 62.10 66.00 68.99 50.37 51.81 55.74 59.77 44.36 45.12 49.00 54.45 % FP LASSO 4.00 10.10 39.78 41.37 1.38 4.23 9.71 7.96 0.70 2.54 5.84 4.97 10.76 14.20 41.74 44.22 5.34 7.07 18.36 17.43 3.16 4.41 10.10 9.54 LASSOy 3.88 9.73 39.60 41.21 1.36 4.18 9.30 7.31 0.68 2.54 5.58 4.63 11.01 12.54 21.19 25.62 5.31 6.18 9.89 10.63 3.16 3.97 5.91 6.19 GLS-LAS 3.29 4.00 22.44 27.49 1.11 1.38 6.71 6.94 0.52 0.77 4.29 4.56 10.47 10.96 22.42 28.41 4.94 5.27 12.68 12.75 2.96 3.34 7.62 7.56 ARDL-LAS 1.93 4.38 21.00 21.19 0.69 2.24 5.60 4.55 0.35 1.30 3.30 2.80 5.03 7.59 13.53 14.99 2.57 3.56 5.63 5.98 3.25 2.39 3.14 3.27 FaSel 23.74 29.11 43.43 45.20 7.05 8.36 12.49 10.81 3.83 3.86 6.80 6.20 0.89 5.37 41.61 43.83 0.25 2.41 19.71 20.08 46.98 29.29 11.30 11.42 ARMAr-LAS 4.74 4.74 5.06 4.88 1.68 1.46 1.41 1.05 0.80 0.79 0.79 0.66 12.00 12.56 12.64 12.79 5.81 6.12 6.36 6.46 3.55 3.78 3.91 4.10 10 CoEr LASSOy 1.00 0.99 0.98 0.98 1.00 1.00 0.99 0.99 1.00 1.00 0.99 1.00 1.00 0.98 0.93 0.94 1.00 0.98 0.97 0.97 1.00 0.99 0.98 0.97 GLS-LAS 0.93 0.75 0.76 0.81 0.94 0.77 0.77 0.72 0.93 0.79 0.83 0.75 0.93 0.77 0.77 0.81 0.95 0.81 0.87 0.83 0.95 0.84 0.90 0.86 ARDL-LAS 1.01 0.98 0.84 0.86 1.01 0.99 0.96 0.97 1.01 0.99 0.96 0.99 0.99 0.94 0.73 0.74 0.99 0.95 0.88 0.92 0.99 0.96 0.92 0.94 FaSel 3.80 2.78 1.54 1.35 2.14 1.42 0.99 0.89 1.30 1.11 0.93 0.85 1.24 1.17 1.09 1.09 1.16 1.10 1.07 1.02 1.50 1.26 1.04 0.99 ARMAr-LAS 0.96 0.74 0.45 0.44 0.96 0.76 0.52 0.47 0.97 0.77 0.56 0.48 0.94 0.73 0.46 0.44 0.95 0.75 0.58 0.54 0.95 0.78 0.62 0.58 RMSFE LASSOy 1.00 0.99 0.99 0.99 1.00 0.99 0.98 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.94 0.96 1.00 0.99 0.98 0.98 1.00 0.99 0.98 0.97 GLS-LAS 0.96 0.83 0.82 0.83 0.94 0.83 0.79 0.69 0.95 0.85 0.81 0.73 0.96 0.84 0.83 0.84 0.96 0.86 0.84 0.78 0.97 0.87 0.88 0.81 ARDL-LAS 1.01 1.01 1.01 1.01 1.01 1.01 0.99 0.99 1.01 1.02 1.01 1.01 1.01 0.98 0.88 0.90 1.01 0.98 0.92 0.93 1.01 0.99 0.96 0.94 FaSel 1.06 0.98 0.97 0.97 1.03 0.99 0.91 0.87 0.97 0.96 0.90 0.86 1.05 1.03 0.99 0.98 1.05 1.03 0.90 0.88 1.08 1.01 0.93 0.85 ARMAr-LAS 0.98 0.83 0.66 0.65 0.99 0.85 0.67 0.59 1.01 0.85 0.69 0.60 0.97 0.81 0.67 0.65 0.96 0.85 0.70 0.64 0.98 0.84 0.75 0.68 % TP LASSO 100.00 100.00 99.55 99.44 100.00 99.97 99.28 98.40 100.00 99.99 99.62 98.57 98.58 95.16 90.73 92.28 97.32 93.40 87.05 85.58 96.21 91.83 85.90 82.66 LASSOy 100.00 99.99 99.62 99.49 100.00 99.97 99.26 98.31 100.00 99.99 99.64 98.43 98.59 95.38 90.25 91.82 97.30 93.47 85.98 84.76 96.24 91.99 85.12 82.30 GLS-LAS 100.00 100.00 99.88 99.70 100.00 100.00 99.87 99.75 100.00 100.00 99.86 99.79 99.03 98.69 93.08 93.32 97.88 97.41 90.86 91.45 97.15 96.03 89.02 89.01 ARDL-LAS 100.00 100.00 99.50 99.43 100.00 99.97 99.33 98.05 100.00 99.99 99.67 98.28 98.48 95.50 91.52 92.47 97.20 93.33 87.15 85.91 96.16 91.81 85.29 82.94 FaSel 56.65 68.47 86.44 92.30 86.67 95.87 99.41 99.25 99.13 99.70 99.81 99.35 91.30 87.32 85.34 88.19 89.63 86.16 89.02 89.03 89.95 86.49 87.42 87.79 ARMAr-LAS 100.00 100.00 100.00 100.00 100.00 99.99 100.00 100.00 100.00 100.00 100.00 100.00 98.98 99.07 99.04 99.46 98.03 98.35 98.40 98.80 97.32 97.34 97.93 97.98 % FP LASSO 4.01 10.46 39.21 40.38 1.33 4.22 9.53 7.87 0.72 2.65 5.95 5.02 14.51 18.97 43.24 45.34 7.37 9.94 14.84 12.14 4.71 6.58 8.73 7.21 LASSOy 3.83 10.11 38.64 40.45 1.29 4.20 9.07 7.31 0.71 2.62 5.75 4.69 14.22 18.51 40.75 42.73 7.27 9.58 13.68 11.25 4.70 6.28 8.17 6.79 GLS-LAS 3.40 4.13 21.83 26.86 1.03 1.37 6.52 7.00 0.54 0.83 4.57 4.60 13.71 14.25 28.11 32.01 7.07 7.66 12.24 11.48 4.53 5.17 7.60 6.99 ARDL-LAS 1.98 4.71 20.30 21.49 0.65 2.20 5.56 4.53 0.35 1.48 3.48 2.83 6.48 9.56 20.53 21.93 3.54 4.88 7.20 6.02 2.31 3.17 4.22 3.64 FaSel 23.08 30.52 43.88 44.27 7.16 8.81 12.50 10.90 3.15 3.84 6.88 6.21 10.95 21.29 46.16 49.11 2.70 8.69 22.00 20.23 12.72 13.11 12.77 12.07 ARMAr-LAS 4.66 4.60 4.61 4.94 1.53 1.56 1.42 1.04 0.82 0.76 0.74 0.64 15.00 15.12 15.44 15.32 7.78 7.85 8.02 7.72 5.05 5.07 4.99 4.99
@@
Table 2: DGPs (C) and (D). CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO. For each $n$ setting the best CoEr and RMSFE are in bold.
SNR 1 10
DGP (C) (D) (C) (D)
n 50 150 300 50 150 300 50 150 300 50 150 300
@@
LASSOy 0.52 0.65 0.74 0.51 0.63 0.69 1.05 1.12 1.15 1.03 1.10 1.11
GLS-LAS 0.65 0.88 0.93 0.63 0.87 0.91 0.83 0.92 0.97 0.82 0.92 0.97
ARDL-LAS 0.44 0.56 0.64 0.44 0.54 0.60 0.78 0.89 0.91 0.77 0.88 0.91
FaSel 1.07 1.04 1.10 1.02 1.04 1.05 1.26 1.07 1.03 1.09 1.06 1.03
ARMAr-LAS 0.50 0.61 0.67 0.48 0.56 0.60 0.70 0.83 0.86 0.65 0.76 0.80
@@
LASSOy 0.78 0.86 0.93 0.79 0.84 0.86 0.94 0.98 0.99 0.96 0.96 0.98
GLS-LAS 0.78 0.90 0.94 0.79 0.90 0.93 0.81 0.90 0.95 0.83 0.89 0.95
ARDL-LAS 0.75 0.86 0.94 0.75 0.80 0.83 0.93 0.97 0.99 0.91 0.92 0.95
FaSel 0.97 0.95 0.93 0.99 0.99 0.98 0.96 0.94 0.93 0.96 0.94 0.93
ARMAr-LAS 0.64 0.72 0.77 0.66 0.70 0.74 0.66 0.74 0.77 0.66 0.67 0.73
@@
LASSO 61.60 55.70 55.00 55.70 43.30 36.00 88.70 89.20 89.90 85.90 83.10 80.00
LASSOy 50.20 48.60 48.40 38.20 30.80 26.60 88.70 88.90 89.70 84.90 82.50 79.50
GLS-LAS 58.40 55.30 53.90 48.60 41.50 35.20 90.30 88.70 89.70 86.40 84.50 80.90
ARDL-LAS 52.50 47.50 46.20 42.10 33.20 27.50 87.80 88.60 89.40 84.90 82.20 79.70
FaSel 49.30 50.00 52.60 50.60 41.10 35.40 68.20 85.40 89.30 80.90 84.30 82.40
ARMAr-LAS 70.30 67.50 65.60 62.60 55.40 47.90 98.80 98.50 98.10 98.40 97.80 96.70
@@
LASSO 37.20 18.90 11.00 39.30 20.50 11.60 35.00 14.50 9.10 40.00 16.40 9.90
LASSOy 15.40 7.40 4.80 17.10 8.80 5.40 33.90 13.90 8.70 37.40 15.20 9.40
GLS-LAS 17.10 14.40 9.00 19.50 16.20 9.90 20.00 10.50 7.80 26.70 13.90 9.20
ARDL-LAS 6.40 2.80 1.80 7.30 3.40 2.00 10.90 5.70 3.50 11.40 5.60 3.40
FaSel 35.40 18.60 11.80 38.50 21.80 13.00 39.30 17.20 9.90 45.40 23.20 13.90
ARMAr-LAS 6.70 2.00 1.00 14.40 6.40 3.90 7.90 2.50 1.20 17.20 8.20 5.30
@@ 4.2 Empirical Application
We consider Euro Area (EA) data composed of 309 monthly macroeconomic time series spanning the period between January 1997 and December 2018. The series are listed in Supplement I, grouped according to their measurement domain: Industry & Construction Survey (ICS), Consumer Confidence Indicators (CCI), Money & Interest Rates (M&IR), Industrial Production (IP), Harmonized Consumer Price Index (HCPI), Producer Price Index (PPI), Turnover & Retail Sale (TO), Harmonized Unemployment Rate (HUR), and Service Surveys (SI). Supplement I also reports transformations applied to the series to achieve stationarity (we did not attempt to identify or remove outliers), as well as an analysis of the autocorrelation functions that justifies the use of our ARMAr-LASSO in this context. The target variable is the Overall EA Consumer Price Index (CPI), which is transformed as I(2) (i.e. integration of order 2) following Stock and Watson (2002b):
$y_{t+h}=(1200/h)\text{log}(CPI_{t+h}/CPI_{t})-1200\ \text{log}(CPI_{t}/CPI_{t-1})\ \ ,$
where $y_{t}=1200\ \text{log}(CPI_{t}/CPI_{t-1})-1200\ \text{log}(CPI_{t-1}/CPI_{t-2})$, and $h$ is the forecasting horizon. We compute forecasts of $y_{t+h}$ at horizons $h=12$ and $24$ using a rolling $\omega$-year window $[t-\omega,t+1]$; the models are re-estimated at each $t$, adding one observation on the right of the window and removing one observation on the left. The last forecast is December 2018. The methods employed for our empirical exercise are:
@@
    LAS (Tibshirani, 1996): forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=0}^{11}\widehat{\phi}_{l}y_{t-l}+\sum_{i=1}^{308}\widehat{\alpha}_{i}x_{it},$ where $(\widehat{\phi}_{0},\dots,\widehat{\phi}_{11},\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO.
@@
    GLS-LAS (Chronopoulos et al., 2023): forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=1}^{p_{\varepsilon}}\widehat{\phi}_{l}y_{t-l+1}+\sum_{i=1}^{308}\widehat{\alpha}_{i}\widetilde{x}_{it},$ where $(\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO using pre-filtered response and predictors (the $\widetilde{x}$’s) as detailed in Supplement D.
@@
    FaSel (Fan et al., 2020): FarmSelector applies feature selection on factors residuals. Forecasts are obtained from the equation: $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\widehat{\boldsymbol{\Lambda}}\widehat{\boldsymbol{f}}_{t}+\widehat{\boldsymbol{\alpha}}^{\prime}\widehat{\mathbf{z}}_{t}+\sum_{i=1}^{p}\widehat{\phi}_{i}y_{t-i+1},$ where $\widehat{\boldsymbol{f}}_{t}$ is a $r$-dimensional vector of factors estimated with PCA (as in Stock and Watson (2002a, b)), $\widehat{\mathbf{z}}_{t}=\widehat{\boldsymbol{\Lambda}}\widehat{\boldsymbol{f}}_{t}-\mathbf{x}_{t}$, $\widehat{\boldsymbol{\Lambda}}^{\prime}$ is the $n\times r$ matrix of loadings, and $\widehat{\boldsymbol{\alpha}}$ is the sparse vector obtained applying the LASSO. The number of factors $r$ is chosen with the approach described in Ahn and Horenstein (2013).
@@
Table 3 reports ratios of RMSFEs between pairs of methods (RMSFE (ratio)), as well as significance of the corresponding Diebold-Mariano test (Diebold and Mariano, 1995). Furthermore, the column “Selected Variables (Av.)” reports the average number of selected variables with ARMAr-LAS (on the left side of the vertical bar), and other LASSO-based methods (on the right side of the vertical bar). Notably, ARMAr-LAS produces significantly better forecasts than AR($p$) and LASSO-based methods, and provides a more parsimonious model than the LAS, ARDL-LAS and FaSel. This is, in principle, consistent with the theoretical analysis we provided earlier. The sparser ARMAr-LAS output may be due to fewer false positives, as compared to other methods which suffer from the effects of spurious correlations induced by serial correlation. Notably, GLS-LAS selects fewer predictors than ARMAr-LAS but provides significantly worse predictions. However, since in this real data application we do not know the true DGP, any comment regarding accuracy in feature selection is necessarily speculative.

### v3

@@ 4.1 Simulation Experiments
Table 1: DGP (A). CoEr, RMSFE (relative to WM), %TP and %FP for various ARMAr-LASSO settings. For each $T$ setting the best CoEr and RMSFE are in bold.
SNR = 1 $T$ 75 150 300 CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP ARMAr-LAS 1.06 1.02 0.49 0.13 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 $y_{t-1}$ 1.30 1.20 0.49 0.15 1.08 1.21 0.61 0.01 1.08 1.24 0.77 0.01 $y_{t-1}^{(3)}$ 1.06 1.02 0.44 0.18 1.02 1.00 0.52 0.01 1.00 1.00 0.69 0.01 AR(1) 1.02 1.00 0.49 0.16 1.03 1.05 0.61 0.01 1.02 1.06 0.75 0.01 AR(2) 1.10 1.02 0.48 0.15 1.02 1.00 0.57 0.02 1.00 1.00 0.72 0.01 AR(3) 1.09 1.00 0.49 0.15 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 AIC 1.06 1.01 0.49 0.15 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 AICc 1.05 0.99 0.49 0.14 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 blCV 1.07 1.00 0.49 0.14 1.02 1.01 0.61 0.01 1.02 1.01 0.77 0.01 SNR = 10 $T$ 75 150 300 CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP ARMAr-LAS 1.05 1.03 0.87 0.05 1.03 1.01 0.97 0.02 1.00 1.00 1.00 0.01 $y_{t-1}$ 1.69 1.91 0.86 0.05 1.63 1.91 0.97 0.02 1.81 1.86 1.00 0.01 $y_{t-1}^{(3)}$ 1.05 1.03 0.65 0.09 1.03 1.01 0.80 0.01 1.00 1.00 0.93 0.01 AR(1) 1.13 1.04 0.85 0.06 1.17 1.12 0.96 0.02 1.19 1.10 1.00 0.01 AR(2) 1.05 1.01 0.84 0.07 1.03 1.01 0.95 0.03 1.00 1.00 0.99 0.02 AR(3) 1.07 1.01 0.86 0.05 1.03 1.01 0.97 0.02 1.00 1.00 1.00 0.01 AIC 1.06 1.02 0.86 0.05 1.03 1.02 0.97 0.02 1.00 1.00 1.00 0.01 AICc 1.05 1.02 0.85 0.05 1.03 1.02 0.97 0.02 1.00 1.00 1.00 0.01 blCV 1.07 1.04 0.85 0.05 1.10 1.09 0.97 0.02 1.15 1.07 1.00 0.01
@@
Table 2: DGP (B). CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO. For each $n$ setting the best CoEr and RMSFE are in bold.
$q=0$ $q=1$
SNR 1 10 1 10
n 75 150 300 75 150 300 75 150 300 75 150 300
@@
LASSOy 0.49 0.59 0.66 0.92 0.95 0.97 0.50 0.56 0.63 0.93 0.95 0.97
GLS-LAS 0.76 0.87 0.92 0.90 0.93 0.97 0.72 0.85 0.91 0.90 0.93 0.97
ARDL-LAS 0.45 0.54 0.61 0.81 0.89 0.91 0.46 0.52 0.58 0.80 0.87 0.91
FaSel 1.04 1.03 1.06 1.17 1.07 1.01 1.02 1.04 1.04 1.06 1.06 1.03
ARMAr-LAS 0.41 0.50 0.55 0.53 0.60 0.64 0.43 0.48 0.53 0.52 0.59 0.62
@@
LASSOy 0.80 0.88 0.90 0.97 0.97 0.98 0.81 0.83 0.87 0.94 0.97 0.98
GLS-LAS 0.82 0.88 0.94 0.88 0.91 0.96 0.82 0.89 0.92 0.88 0.91 0.95
ARDL-LAS 0.80 0.89 0.90 1.00 0.99 0.97 0.78 0.80 0.85 0.92 0.93 0.98
FaSel 0.96 0.95 0.96 0.96 0.94 0.94 0.99 0.97 0.98 0.98 0.94 0.96
ARMAr-LAS 0.67 0.74 0.79 0.71 0.75 0.80 0.70 0.74 0.79 0.70 0.74 0.79
@@
LASSO 57.10 52.00 52.60 87.60 87.00 87.50 48.10 39.40 33.40 82.10 79.30 76.60
LASSOy 46.50 43.80 44.10 87.20 86.50 87.40 31.90 27.60 22.90 81.70 78.40 76.00
GLS-LAS 55.30 51.10 51.30 89.10 86.30 86.20 43.30 37.90 32.40 83.30 80.70 77.30
ARDL-LAS 46.70 42.20 42.20 85.90 86.50 87.20 35.00 29.70 24.00 81.10 78.60 75.20
FaSel 47.10 45.60 49.70 73.60 81.80 87.10 45.60 37.90 32.10 78.60 79.00 78.20
ARMAr-LAS 65.60 62.40 60.80 97.20 96.60 96.40 54.80 50.00 43.40 96.40 95.30 94.30
@@
LASSO 32.30 18.40 11.30 29.50 14.60 9.10 31.80 19.70 11.30 31.50 16.50 9.70
LASSOy 10.90 6.50 4.40 27.60 13.50 8.70 12.20 7.80 5.00 29.50 15.30 9.20
GLS-LAS 19.70 13.80 9.30 21.60 10.50 7.70 19.60 15.40 9.80 25.30 14.20 9.10
ARDL-LAS 4.10 2.30 1.70 8.60 5.70 3.60 5.20 3.10 1.80 8.70 5.40 3.20
FaSel 29.80 17.70 11.60 32.20 16.50 9.60 32.30 21.10 12.70 35.90 22.40 13.50
ARMAr-LAS 4.10 2.00 0.90 5.10 2.40 1.20 10.50 6.20 3.60 13.30 8.30 5.10
In addition to the results presented in this section, Supplement D reports further simulations based on a common AR(1) structure with varying autoregressive coefficients $\phi$, larger sample sizes $T$, and cases where the ARMAr-LASSO misspecifies the AR model of the predictors.
@@ 4.2 Empirical Application
We consider Euro Area (EA) data composed of 309 monthly macroeconomic time series spanning the period between January 1997 and December 2018. We note that the data used here are not real-time vintages. Rather, all macroeconomic series were downloaded at a single point in time and correspond to the most recent available releases as of the download date. The series are listed in Supplement E, grouped according to their measurement domain: Industry & Construction Survey (ICS), Consumer Confidence Indicators (CCI), Money & Interest Rates (M&IR), Industrial Production (IP), Harmonized Consumer Price Index (HCPI), Producer Price Index (PPI), Turnover & Retail Sale (TO), Harmonized Unemployment Rate (HUR), and Service Surveys (SI). Supplement E also reports transformations applied to the series to achieve stationarity (we did not attempt to identify or remove outliers), as well as an analysis of the autocorrelation functions that justifies the use of our ARMAr-LASSO in this context. The target variable is the Overall EA Consumer Price Index (CPI), which is transformed as I(2) (i.e. integration of order 2) following SeW2002b: $y_{t+h}=(1200/h)\text{log}(CPI_{t+h}/CPI_{t})-1200\ \text{log}(CPI_{t}/CPI_{t-1})$, where $y_{t}=1200\ \text{log}(CPI_{t}/CPI_{t-1})-1200\ \text{log}(CPI_{t-1}/CPI_{t-2})$, and $h$ is the forecasting horizon. We compute forecasts of $y_{t+h}$ at horizons $h=12$ and $24$ using a rolling $\omega$-year window $t-\omega,t+1]$; the models are re-estimated at each $t$, adding one observation on the right of the window and removing one observation on the left. The last forecast is December 2018. The methods employed for our empirical exercise are:
@@
    LAS (tibshirani96): forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=0}^{11}\widehat{\phi}_{l}y_{t-l}+\sum_{i=1}^{308}\widehat{\alpha}_{i}x_{it},$ where $(\widehat{\phi}_{0},\dots,\widehat{\phi}_{11},\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO.
@@
    GLS-LAS (chronopoulos2023): forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=1}^{p_{\varepsilon}}\widehat{\phi}_{l}y_{t-l+1}+\sum_{i=1}^{308}\widehat{\alpha}_{i}\widetilde{x}_{it},$ where $(\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO using pre-filtered response and predictors (the $\widetilde{x}$’s) as detailed in Supplement C.
@@
    FaSel (Fan2020): FarmSelector applies feature selection on factors residuals. Forecasts are obtained from the equation: $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\widehat{\boldsymbol{\Lambda}}\widehat{\boldsymbol{f}}_{t}+\widehat{\boldsymbol{\alpha}}^{\prime}\widehat{\mathbf{z}}_{t}+\sum_{i=1}^{p}\widehat{\phi}_{i}y_{t-i+1},$ where $\widehat{\boldsymbol{f}}_{t}$ is a $r$-dimensional vector of factors estimated with PCA (as in SeW2002a; SeW2002b), $\widehat{\mathbf{z}}_{t}=\widehat{\boldsymbol{\Lambda}}\widehat{\boldsymbol{f}}_{t}-\mathbf{x}_{t}$, $\widehat{\boldsymbol{\Lambda}}^{\prime}$ is the $n\times r$ matrix of loadings, and $\widehat{\boldsymbol{\alpha}}$ is the sparse vector obtained applying the LASSO. The number of factors $r$ is chosen with the approach described in AhnHorenstein2013.
@@
Table 3 reports ratios of RMSFEs between pairs of methods (RMSFE (ratio)), as well as significance of the corresponding Diebold-Mariano test (DiebMar1995). Furthermore, the column “Selected Variables (Av.)” reports the average number of selected variables with ARMAr-LAS (on the left side of the vertical bar), and other LASSO-based methods (on the right side of the vertical bar). Notably, ARMAr-LAS produces significantly better forecasts than AR($p$) and LASSO-based methods, and provides a more parsimonious model than the LAS, ARDL-LAS and FaSel. This is, in principle, consistent with the theoretical analysis we provided earlier. The sparser ARMAr-LAS output may be due to fewer false positives, as compared to other methods which suffer from the effects of spurious correlations induced by serial correlation. Notably, GLS-LAS selects fewer predictors than ARMAr-LAS but provides significantly worse predictions. However, since in this real data application we do not know the true DGP, any comment regarding accuracy in feature selection is necessarily speculative.
Table 3: RMSFE (ratio): ratios of RMSFE contrasting pairs of employed methods; for each ratio, we perform a Diebold-Mariano test (alternative: the second method is less accurate in forecasting) and report p-values as 0 ’***’ 0.001 ’**’ 0.01 ’*’ $^{\mathbin{\vbox{\hbox{\scalebox{0.5}{$\bullet$}}}}}’ 0$.1”. Selected Variables (Av.): average of the number of variables selected by ARMAr-LAS (left side of the vertical bar) and LASSO-based benchmarks (right side of the vertical bar).
Method 1 Method 2 RMSFE (ratio) Selected Variables (Av.)
hh=12 $h=24$ hh=12 $h=24$
ARMAr-LAS LAS 0.69*** 0.82* 6.0||67.9 6.2||60.9
ARMAr-LAS GLS-LAS 0.66*** 0.61*** 6.0||3.5 6.2||3.8